# Knowledge Graph Based Agent for Complex, Knowledge-Intensive QA in Medicine

### v1

@@ 3.3 Revise and Answer Actions
If $F$ has triplets, KGARevion calls the Revise action to adjust the triplets in $F$ to include more knowledge that helps with the answering of the input question. The head and tail entities of the revised triplets are then reviewed by the Review action to make sure that they are correct and related to the input question. If the Review action outputs “True”, then the revised triplets are added to the set of True triplets $V$. Otherwise, KGARevion continues to call the Revise action until the max round $k$ ($k$geq 1) is achieved.
After obtaining the set of True triplets from the Review or Revise actions, KGARevion finally calls the Answer action to prompt the LLM to select the most suitable answer $y$ from the set of answer candidates $C$ of input question $q$ based on the triplets in $V$, where $y=P(q,V,C)$ and $y\in C$.
@@ 4 Results
Datasets. We first start with four multi-choice medical QA benchmarks (Xiong et al., 2024a) (Table 1). In addition, we introduce a new benchmark for multi-choice complex medical QA focused on differential diagnosis (DDx), named MedDDx. We begin by collecting questions and corresponding answers from STaRK-Prime (Wu et al., 2024c). For each question, we then select the top three entities with the highest semantic similarity to serve as additional answer candidates. MedDDx comprises a total of 1,769 multi-choice QA samples. Based on the standard deviation of semantic similarity between answer candidates and the correct answer, we categorize the dataset into three difficulty levels: MedDDx-Basic, MedDDx-Intermediate, and MedDDx-Expert (The samples in each dataset are shown in Fig. 2b, and details are available in Appendix 5).
@@
Baselines. We consider 8 LLM-based reasoning models, 4 RAG-based models, and 3 KG-based models. The LLM-based reasoning models include LLaMA (2-7B/13B, 3-8B, 3.1-8B) (Touvron et al., 2023; Dubey et al., 2024), Mistral (Jiang et al., 2023), MedAlpaca (7B) (Han et al., 2023), PMC-LLaMA (7B) (Wu et al., 2024a), LLaMA3-OpenBioLLM-8B (Ankit Pal, 2024), and MEDITRON (Chen et al., 2023). The RAG-based models include Self-RAG (Asai et al., 2024), MedRAG (Xiong et al., 2024b), KG-RAG (Soman et al., 2023), and KG-Rank (Yang et al., 2024). The KG-based models include QAGNN (Yasunaga et al., 2021), JointLK (Sun et al., 2022), and Dragon (Yasunaga et al., 2022).
@@
Method MMLU-Med MedQA-US PubMedQA* BioASQ-Y/N Basic Intermidate Expert
@@
4.1 Benchmarking KGARevion under Multi-choice Reasoning Setting
@@
4.2 Benchmarking KGARevion under Open-ended Reasoning Setting
We transform multiple-choice questions into descriptive, open-ended ones to better simulate real-world medical scenarios, where such inquiries are more common (see details in Appendix A.3). This adjustment requires our model to generate responses without predefined choices, encouraging holistic reasoning and the integration of diverse knowledge sources. By removing answer choices, we can more effectively assess the reasoning ability of KGARevion in complex medical situations, resulting in a more realistic evaluation of its capabilities. Table 3 shows the accuracy and variance across all datasets. The variance here denotes the difference in accuracy compared with that in the multi-choice reasoning setting.
@@ 4.3 Ablation Analyses
Effect of the ‘Review’ action. As shown in Table 2, 3, and Fig. 4 (KGARevion (w/o Review) vs. KGARevion (w/o Revise))), the Review action plays an important role in answering medical questions under two settings, which improves the average accuracy across all datasets by 3.3% and 3%, respectively. Fig. 4 shows that the Review action has a more pronounced effect on the MedDDx dataset than the four gold-standard datasets under two settings, suggesting that its integration enhances the model’s ability to tackle complex medical questions. Furthermore, the Review action leads to greater accuracy improvements in the four gold-standard datasets under the open-ended reasoning setting compared to the multiple-choice reasoning setting. This highlights the significance of verifying generated answers, particularly in an open-ended reasoning setup.
@@
Number of refinement rounds in the ‘Revise’ action. The Revise action is designed to enhance accuracy by correcting erroneous triplets until they are verified as true by the Review action. Tables 2 and 3, and Fig. 4 demonstrate its positive impact on KGARevion across both settings. Specifically, Figure 4 indicates that the Review action significantly improves performance on MedDDx dataset, yielding average enhancements of 9% and 4% in accuracy for both settings compared to questions in the gold-standard datasets. Additionally, we investigate the impact of the number of revision rounds across all datasets in both settings, as shown in Table 2 and 3. The results indicate that KGARevion can achieve optimal performance with $k=1$ on most of datasets in the multi-choice reasoning setting. However, it benefits from additional iterations when addressing complex questions, such as those in the MedDDx-Expert dataset. In the open-ended reasoning setting, KGARevion typically requires more iterations to arrive at the correct answer.
@@ 4.4 Versatility of KGARevion
KGARevion can be used with different LLMs. KGARevion is a versatile agent that can be implemented with a variety of LLMs. We implement KGARevion using three distinct models: LLaMA3-8B, LLaMA3.1-8B, and GPT-4o. The averaged results across all datasets, as shown in Fig. 5a, demonstrate the effectiveness of KGARevion’s architecture, consistently improving the performance of the backbone LLMs by 6%, 7%, and 2%, respectively.
@@
KGARevion can be used with different medical KGs. The Review action in KGARevion grounds the generated triplets using KGs. To evaluate the impact of different KGs (see details in Appendix B.5), we implement KGARevion with two comprehensive KGs and assess its performance across all datasets, as shown in Fig. 5b. The results show that KGARevion is not sensitive to the choice of knowledge bases, highlighting its robustness and generalizability, despite PrimeKG (Chandak et al., 2023) being much larger than OGB-biokg (Hu et al., 2020). This robustness arises because KGs are used only in the Review action to verify generated triplets rather than as a source for retrieving knowledge. This also explains why KGARevion outperforms KG-based RAG models, which heavily rely on the chosen KGs, whereas KGARevion uses comprehensive KGs simply to ensure that the generated triplets are aligned with medical knowledge.
@@ 4.5 Sensitivity Analyses
Recent studies have revealed that LLMs can be surprisingly sensitive to both how candidate answers are ordered and indexed in multi-choice setups (Zheng et al., 2023; Pezeshkpour & Hruschka, 2023). These studies found that LLMs are not robust multiple-choice selectors and exhibit order sensitivity, favoring answers at the first position (Li et al., 2024). To investigate this issue, we examine how the order and indexing of answers affect model performance. We evaluate KGARevion using LLaMA3-8B and LLaMA3.1-8B as the backbone and compare its performance with their LLM-only counterparts across all datasets (details are provided in Appendix B.4.1). Fig. 6 illustrates the changes in accuracy when the order or labels of the candidate answers are altered.
Ordering of candidate answers in multi-choice setups. Fig. 6a shows that pure LLMs are sensitive to answer order, with an average accuracy shift of 8.4% for LLaMA3-8B and 16.0% for LLaMA3.1-8B. In contrast, KGARevion demonstrates significantly greater robustness to answer order. This robustness is primarily due to KGARevion’s ability to fairly evaluate each answer using the Generate action, effectively mitigating the impact of answer order on model performance.
@@
Indexing of candidate answers in multi-choice setups. The accuracy of pure LLMs shows a substantial shift when relabeling answers from ABCD to EFGH, as illustrated in Fig. 6b. Specifically, the average accuracy shift is 8.1% for LLaMA3-8B and 12.9% for LLaMA3.1-8B. In contrast, our agent KGARevion significantly improves the stability of these LLMs, reducing the accuracy loss to 2.59% and 3.86%, respectively. This finding further highlights the robustness of KGARevion.
Figure 7: The reasoning process of KGARevion under multi-choice setting (Left) and open reasoning setting (Right).
4.6 Case Studies
Fig. 7 illustrates the reasoning process of KGARevion in both settings, using the same input question. In both cases, KGARevion arrives at the correct answer, but the reasoning processes differ. In the open reasoning setting, KGARevion requires more iterations to revise triplets and guide the reasoning compared to the multiple-choice setting. Additionally, the verified correct triplets provide a reasoning path that helps explain the final answer, such as ’partial deletion of the long arm of chromosome 19 $\rightarrow$ associated with $\rightarrow$ 19q13.11 deletion syndrome and partial chromosome 19 deletions $\rightarrow$ associated with $\rightarrow$ disease’.
@@ 5 Conclusion
Medical reasoning presents unique challenges that require integrating multi-source, grounded, and specialized domain knowledge. In this work, we introduced KGARevion, a KG-based LLM agent that addresses these challenges by combining the non-codified knowledge of LLMs with the structured, codified knowledge of medical concepts stored in KGs. Through its adaptive reasoning and mechanisms for generating, verifying, and revising knowledge, KGARevion can handle complex medical QA. Experiments across multiple-choice and open-ended tasks, using a variety of datasets—including challenging new benchmarks—demonstrate KGARevion’s ability to systematically improve accuracy. By grounding LLM-generated knowledge in KGs, KGARevion ensures contextual relevance and reliability, making it a valuable tool for knowledge-intensive medical QA.
6 Acknowledgement
We gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, ARPA-H BDF Toolbox Program, awards from Pfizer Research, Harvard Data Science Initiative, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH Award, Chan Zuckerberg Initiative, John and Virginia Kaneb Fellowship award at Harvard Medical School, Biswas Computational Biology Initiative in partnership with the Milken Institute, Kempner Institute for the Study of Natural and Artificial Intelligence, and Harvard Medical School Dean’s Innovation Awards. V.G. is supported by the UK Medical Research Council, MR/W00710X/1. The content is solely the responsibility of the authors.
@@ 6 Acknowledgement
Appendix A Datasets
A.1 Gold Standard Multi-choice Medical QA Dataset
@@
A.2 Newly Created Medical QA Dataset – MedDDx
@@
  • While large language models (LLMs) can perform QA tasks, they often rely heavily on semantic dependencies, making it difficult for them to identify the correct answer among semantically similar answer candidates;
@@
A.3 Conversion of Multi-Choice Type Questions to Descriptive Type
@@
Appendix B Implementation Details
B.1 Experiment Environments
@@
B.2 Fine-tuning Details
@@
B.3 Implementation Details of Baseline Models

### v2

@@ 3.3 Revise and Answer Actions
If $F$ has triplets, KGARevion calls the Revise action to adjust the triplets in $F$ to include more triplets covering more medical concepts that help with the answering of the input question. The new generated revised triplets are then reviewed by the Review action to make sure that they are correct and related to the input question. If the Review action outputs “True”, then the revised triplets are added to the set of True triplets $V$. Otherwise, KGARevion continues to call the Revise action until the max round $k$ ($k$geq 1) is achieved.
After obtaining the set of True triplets from the Review or Revise actions, KGARevion finally calls the Answer action to prompt the LLM to generate the answer to input question $q$ based on the verified triplets in $V$.
@@ 4 Results
Datasets. We first start with four multi-choice medical QA benchmarks (Xiong et al., 2024a) (Table 1). In addition, we introduce a new benchmark for multi-choice complex medical QA focused on differential diagnosis (DDx), named MedDDx. We begin by collecting questions and corresponding answers from STaRK-Prime (Wu et al., 2024c). For each question, we then select the top three entities with the highest semantic similarity to serve as additional answer candidates. MedDDx comprises a total of 1,769 multi-choice QA samples. Based on the standard deviation of semantic similarity between answer candidates and the correct answer, we categorize the dataset into three difficulty levels: MedDDx-Basic, MedDDx-Intermediate, and MedDDx-Expert (The samples in each dataset are shown in Fig. 1, and details are available in Appendix 5).
@@
Baselines. We consider eight LLM-based reasoning models, four RAG-based models, and three KG-based models. The LLM-based reasoning models include LLaMA (2-7B/13B, 3-8B, 3.1-8B) (Touvron et al., 2023; Dubey et al., 2024), Mistral (Jiang et al., 2023), MedAlpaca (7B) (Han et al., 2023), PMC-LLaMA (7B) (Wu et al., 2024a), and LLaMA3-OpenBioLLM-8B (Ankit Pal, 2024). The RAG-based models include Self-RAG (Asai et al., 2024), MedRAG (Xiong et al., 2024b), KG-RAG (Soman et al., 2023), and KG-Rank (Yang et al., 2024). The KG-based models include QAGNN (Yasunaga et al., 2021), JointLK (Sun et al., 2022), and Dragon (Yasunaga et al., 2022).
@@ 4.1 Multi-choice Question-Answering
4.2 Open-Ended Reasoning
We transform multiple-choice questions into descriptive, open-ended ones to better simulate real-world medical scenarios, where such inquiries are more common (see details in Appendix B.4). This adjustment requires our model to generate responses without predefined choices, encouraging holistic reasoning and the integration of diverse knowledge sources. By removing answer choices, we can more effectively assess the reasoning ability of KGARevion in complex medical situations, resulting in a more realistic evaluation of its capabilities. Table 3 shows the accuracy and variance across all datasets. The variance here denotes the difference in accuracy compared with that in the multi-choice reasoning setting.
@@ 4.3 Ablation Analyses
Effect of the ‘Review’ action. As shown in Tables 2 and 3 and in Fig. 4 (KGARevion (w/o Review) vs. KGARevion (w/o Revise)), the Review action plays a crucial role in answering medical questions across both reasoning settings. It improves the average accuracy by 9% in the multiple-choice setting and by 4% in the open-ended setting. Fig. 4 shows that the Review action has a more pronounced effect on the MedDDx dataset than the other four datasets in both reasoning settings. This suggests that incorporating the Review action enhances KGARevion’s performance on complex medical questions. Furthermore, the Review action leads to greater accuracy improvements in the gold-standard datasets under the open-ended reasoning setting than in the multiple-choice setting. These findings emphasize the importance of verifying candidate answers for open-ended reasoning, where retrieval and knowledge validation are critical for accuracy.
@@
Effect of refinement rounds in the ‘Revise’ action. The Revise action enhances accuracy by iteratively correcting erroneous triplets until the Review action verifies them as correct. Tables 2 and 3, along with Fig. 4, demonstrate that KGARevion benefits from this refinement process across both reasoning settings. Fig. 4 shows that the Revise action significantly improves performance on the MedDDx dataset, yielding accuracy gains of 3.3% and 3.0% in the multiple-choice and open-ended settings, respectively. We further analyze the impact of the number of revision rounds across all datasets in both settings, as detailed in Tables 2 and 3. The results indicate that KGARevion achieves optimal performance with $k=1$ in multiple-choice reasoning datasets. However, it benefits from additional rounds of refinement when handling complex questions, such as those in the MedDDx-Expert dataset. KGARevion typically requires more iterations to refine and verify the correct answer for open-ended reasoning, highlighting the importance of iterative knowledge validation in knowledge-intensive medical QA.
@@ 4.4 Versatility of KGARevion
KGARevion integrates with different LLMs. KGARevion is a versatile agent that operates effectively with various LLMs. We implement KGARevion using three models: LLaMA3-8B, LLaMA3.1-8B, and GPT-4-Turbo. The averaged results across all datasets, shown in Fig. 5a, demonstrate that KGARevion consistently enhances the performance of its backbone LLMs, improving accuracy by 6% with LLaMA3-8B, 7% with LLaMA3.1-8B, and 2% with GPT-4-Turbo. These improvements highlight the adaptability of KGARevion across different model architectures, reinforcing its effectiveness in knowledge-intensive QA.
@@
KGARevion adapts to different KGs. The Review action in KGARevion verifies generated triplets using structured knowledge from KGs. To evaluate how different KGs impact performance (see Appendix C.5 for details), we implement KGARevion with two KGs and assess its performance across all datasets, as shown in Fig. 5b. The results show that KGARevion maintains strong performance regardless of the KG used, demonstrating its robustness and generalizability. Despite PrimeKG (Chandak et al., 2023) being significantly larger than OGB-biokg (Hu et al., 2020), KGARevion’s performance remains high.
This stability arises because KGARevion uses KGs exclusively in the Review action to validate generated triplets rather than as a primary knowledge retrieval source. Unlike KG-based RAG models, which depend heavily on the structure and completeness of their KGs, KGARevion only employs KGs to ensure that generated triplets align with medical knowledge. This verification-based approach makes KGARevion more resilient and less dependent on any single knowledge source, allowing it to outperform RAG models that rely extensively on specific KGs for retrieval.
4.5 Evaluation of KGARevion on African Healthcare
To assess whether KGARevion relies on the extensive pre-trained knowledge of LLMs from existing datasets, we evaluate it on AfriMed-QA, a newly published QA dataset released after all baseline models in this study (Olatunji et al., 2024). We benchmark KGARevion and five baselines on this dataset under a multiple-choice setting. Further details on AfriMed-QA are provided in Appendix B.3.
Fig. 6 presents the accuracy of KGARevion compared to the baselines on all expert-level multiple-choice questions in AfriMed-QA. Among the baselines, GPT-4-Turbo achieves the highest accuracy at 65.0%, while LLaMA3.1-8B reaches 55.5%. To demonstrate KGARevion’s effectiveness, we implement it with both LLaMA3.1-8B and GPT-4-Turbo. The results show that KGARevion improves accuracy by 5.2% when using LLaMA3.1-8B and by 4.6% when using GPT-4-Turbo. These findings highlight KGARevion’s robustness on a previously unseen dataset and its strong zero-shot generalization capabilities in underrepresented medical contexts.
Figure 6: Performance of baselines and KGARevion with LLaMA3.1-8B and GPT-4-Turbo on AfriMed-QA expert multichoice questions.
@@ 4.6 Sensitivity Analyses
LLMs often exhibit sensitivity to the ordering and indexing of candidate answers in multiple-choice setups (Zheng et al., 2023; Pezeshkpour & Hruschka, 2023). Prior studies have shown that LLMs are not robust multiple-choice selectors and tend to favor answers appearing in the first position (Li et al., 2024). To examine this issue, we analyze how variations in answer order and indexing impact model performance. We evaluate KGARevion using LLaMA3-8B and LLaMA3.1-8B as backbones and compare its performance with their LLM-only counterparts across all datasets (detailed results in Appendix C.4.1). Fig. 7 presents the accuracy changes when altering the order or relabeling candidate answers.
@@
Impact of answer order in multiple-choice setups. Fig. 7a shows that pure LLMs exhibit strong sensitivity to answer order, with average accuracy shifts of 8.4% for LLaMA3-8B and 16.0% for LLaMA3.1-8B. In contrast, KGARevion demonstrates significantly greater robustness, reducing this sensitivity. This stability arises from KGARevion’s evaluations of answers using its Generate action that mitigates the impact of answer order on performance.
Impact of answer indexing in multiple-choice setups. LLMs also show substantial accuracy shifts when relabeling answers from ABCD to EFGH, as illustrated in Fig. 7b. LLaMA3-8B experiences an average accuracy shift of 8.1%, while LLaMA3.1-8B shows a shift of 12.9%. In contrast, KGARevion significantly stabilizes these LLMs, reducing accuracy loss to 2.59% and 3.86%, respectively. These results highlight KGARevion’s ability to enhance robustness in knowledge-intensive medical reasoning.
@@ 5 Conclusion
Biomedical reasoning requires integrating multi-source, grounded, and specialized domain knowledge to ensure accurate and reliable decision-making. In this work, we introduced KGARevion, a KG-based LLM agent that addresses these challenges by combining the non-codified knowledge embedded in LLMs with the structured, codified knowledge of medical concepts stored in KGs. By dynamically generating, verifying, and refining knowledge, KGARevion adapts to the complexity of biomedical QA to generate accurate and contextually relevant answers. Evaluations across multiple-choice and open-ended tasks, including newly curated challenging benchmarks, demonstrate that KGARevion consistently improves accuracy over existing methods. By grounding LLM-generated knowledge in KGs, KGARevion enhances reasoning robustness and ensures alignment with biomedical knowledge. These results highlight the potential of KGARevion to advance knowledge-intensive medical QA, establishing the foundation for future applications in clinical decision support, biomedical research, and complex medical inference.
Acknowledgment
We gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, Harvard Data Science Initiative, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH, Pfizer Research, Gates Foundation (INV-079038), Chan Zuckerberg Initiative, John and Virginia Kaneb Fellowship at Harvard Medical School, Biswas Computational Biology Initiative in partnership with the Milken Institute, Harvard Medical School Dean’s Innovation Fund for the Use of Artificial Intelligence, and Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. V.G. is supported by the UK Medical Research Council, MR/W00710X/1. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
Ethics Statement
This study evaluates LLMs on benchmark datasets for medical QA. Although LLMs demonstrate promising capabilities in processing and generating medical information, their use in real-world medical decision-making carries inherent risks. LLMs may produce hallucinations (factually incorrect or misleading information), which can lead to inaccurate medical advice. In addition, these models are trained on datasets that can contain biases, which can lead to disparities in healthcare recommendations. Ensuring transparency, interpretability, and fairness in medical AI systems remains an ongoing challenge. Further, medical LLMs should not be considered a substitute for professional medical advice. The deployment of such models must be approached with caution, ensuring alignment with medical ethics, patient safety guidelines, and regulatory standards. Users must be aware of these limitations, and appropriate human oversight is necessary when applying LLMs in clinical or consumer-facing applications.
@@ Acknowledgment
Appendix A Extended Related Work
Fact-checking Models. The vast amount of generated text has created a growing need for fact-checking (Graves & Glaisyer, 2012). Early methods relied on manually defined rules to verify generated statements (Hassan et al., 2014; Wu et al., 2012; Jiang et al., 2011). KGs, as structured and reliable knowledge bases, have also been utilized for fact-checking by analyzing graph connectivity. These methods (Tunstall-Pedoe, 2010; Shi & Weninger, 2016; Ciampaglia et al., 2015) validate generated statements by identifying paths between entities mentioned in the statement. Additionally, some approaches use popular KG representation learning algorithms for KG completion tasks to verify the accuracy of statements (Li et al., 2011). With the rise of LLMs, new fact-checking approaches, e.g. RAG-based models, have emerged to ensure the consistency and reliability of generated outputs by integrating external evidence retrieval for verifying and supporting LLM-generated content (Asai et al., 2024).
Appendix B Datasets
B.1 Gold Standard Multi-choice Medical QA Dataset
@@
B.2 MedDDx
@@
  • While LLMs can perform QA tasks, they often rely heavily on semantic dependencies, making it difficult for them to identify the correct answer among semantically similar answer candidates;
@@
B.3 AfriMed-QA
AfriMed-QA creates a novel multispecialty open-source dataset of 15,275 pan-African clinically diverse QA to rigorously evaluate LLMs in African healthcare. The dataset is sourced from over 500 clinical and non-clinical contributors across 16 countries and covers 32 clinical specialties. First, because the dataset is sourced from healthcare systems whose data are not online, it means that none of the multiple-choice and open-ended questions and answers can be in the knowledgebase of the LLM. Because of that, this benchmark has not been leaked into KGARevion’s LLM. Furthermore, the knowledge cut-off date for KGARevion’s LLM was before AfriMed-QA was released. Second, AfriMed-QA is sourced from clinical, private healthcare systems whereas PrimeKG, STaRK-Prime, and OGB-bioKG are sourced from non-clinical, public, biological data repositories. Furthermore, PrimeKG, STaRK-Prime, and OGB-bioKG were all released several years prior to when the AfriMed-QA benchmark was released. As such, AfriMed-QA has certainly not been leaked into KGARevion’s KG.
B.4 Conversion of Multi-Choice Type Questions to Descriptive Type
@@
Appendix C Implementation Details
C.1 Experiment Environments
@@
C.2 Fine-tuning Details
@@
C.3 Implementation Details of Baseline Models